4) Calculate the pH of a 0. 075 MB a (OH), soln.

calculate the ph
Answer #1

1 mol of Ba(OH)2 gives 2 mol of OH-

So,

[OH-] = 2*[Ba(OH)2]

= 2*0.075 M

= 0.150 M

use:

pOH = -log [OH-]

= -log (0.15)

= 0.8239

use:

PH = 14 - pOH

= 14 - 0.8239

= 13.1761

Answer: 13.18